# Fast Penalized Generalized Estimating Equations for Large Longitudinal Functional Datasets

### v1

@@ A.3 Initial FoSR pffr fit
where ${l}({Y}_{i,j}(s),\mathbb{X}_{i,j};\bm{\theta})$ is the log-likelihood evaluated on a single observation of the outcome ${Y}_{i,j}(s)\in\mathbb{R}$ from cluster $i$, at longitudinal observation $j$, at functional domain point $s$. We denote the model parameter $\bm{\theta}$, $\mathbb{S}$ as the penalty matrix, and $\Lambda_{0}$ as an associated diagonal matrix of smoothing parameters. As discussed in the main text, $\mathbb{X}_{i,j}\in\mathbb{R}^{p}$ is constructed as a product of pre-defined basis functions (e.g. B-splines) and the original covariates, $\bm{x}_{i,j}\in\mathbb{R}^{q}$. Model 13 is equivalent to adopting a correlation structure that assumes all observations $Y_{i,j}(s)$ are mutually independent across $i$, $j$, and $s$.
We fit model 13 with the refund package in R (Goldsmith et al.,, 2024) with the pffr function (Scheipl et al.,, 2015). This calls the mgcv package in R (Wood,, 2017) to fit the model with the gam or bam functions. For example, for a model with two covariates, X1 and X2, we use the following code:
@@
where fam is the exponential dispersion family adopted to construct a pseudo-likelihood based estimating equation in the fGEE, spline.basis is the pre-specified basis spline from the mgcv family (e.g. bs, ps, tp), knots is the number of knots, and m.pffr is the penalty type. For example, as noted in the pffr function, the following
bs.yindex = list(bs="ps", k=5, m=c(2, 1)) indicates 5 cubic B-splines bases with a first order difference penalty.
A.4 Computational Details
Our implementation uses a number of R packages for estimation of nuisance parameters and to improve computational speed. We use the data.table package extensively to increase computational efficiency (Barrett et al.,, 2025). Our code structure is loosely based on the structure from the gee1step) package that implements the One-Step (non-functional) GEE (Lipsitz et al.,, 2017) available on the Github https://github.com/kgoldfeld/gee1step of Professor Keith Goldfeld. We use the Rfast package (Papadakis et al.,, 2018) to estimate the $\rho(s)$ for AR1 correlation and to speed up other standard computations. We use the MASS package to draw multivariate normals (Venables and Ripley,, 2002). We use Allévius, (2018) to estimate and invert covariance matrices with an AR1 structure when the time intervals are irregular. We use the SuperGauss package (Ling and Lysy,, 2022) to invert AR1 matrices with regular time intervals. We use the sanic package to quickly invert other positive definite matrices (Kuschnig,, 2023). As mentioned in the main text and previous Appendix section, we use mgcv (Wood,, 2017) and refund (Goldsmith et al.,, 2024) packages to estimate initial fits and to (optionally) smooth the correlation parameters across the functional domain.
@@ Appendix B Theory
[…]
@@
[…]
@@
[…]
@@
[…]
@@
+$\quad\quad+\left\{\mathbb{P}_{N}\left[\left.\nabla_{\bm{\theta}}\bm{U}_{N}(Z;\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{N}}\right]-\mathbb{E}\left(\left.\nabla_{\bm{\theta}}\bm{U}_{N}(Z;\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{N}}\right)\right\}.$
@@
B.1 Condition (vi)
To provide intuition for condition (vi) in the statement of Theorem B.2, we derive an interpretable set of conditions that imply condition (vi) in a special case of our general framework: a non-functional, univariate (i.e., $p=q=1$), non-clustered (i.e., $n_{i}=1~{}\forall~{}i\in[N]$) ridge regression that is weighted by the inverse of the working covariance matrix. This setting provides insight for the more general case: (1) the fully-iterated unweighted ridge is analogous to the initial penalized GEE estimator that is also unweighted (i.e., adopts an independence working covariance structure) and fully-iterated; (2) the fully-iterated weighted ridge is analogous to the fully-iterated and weighted (i.e., adopts some non-independence working covariance structure) penalized GEE. In this special case, we show that if the scaled smoothing parameters for the initial estimate satisfy $\frac{1}{N}\lambda_{0,N}=O(N^{-1/2})$, and for the weighted estimator satisfy $\frac{1}{N}\lambda_{N}=O(N^{-1/2})$, then condition (vi) holds. That is, if each of these scaled smoothing parameters individually (i.e., no conditions are required jointly on these rates) go to zero fast enough, then $\sqrt{N}\{\bm{M}_{N}(\bm{\theta}_{N})\}^{-1/2}\bm{H}_{N}(\bm{\theta}_{N})(\widehat{\bm{\theta}}^{(0)}_{N}-\bm{\theta}_{N})=O_{\mathbb{P}}(1)$. Importantly, the smoothing parameter rates we require are weaker than (i.e., implied by) the rates needed for the theoretical properties described in Chen et al., (2013). For example, even for the small knot setting, Chen et al., (2013) require that $\lambda=O(N^{\gamma})$ for $\gamma\leq(\tilde{p}+2-\tilde{q})/(2\tilde{p}+3)$. This implies, for instance, that for a $\tilde{p}^{th}$ order truncated polynomial using cubic B-splines, $\gamma\leq(\tilde{p}+2-(\tilde{p}+1))/(2\tilde{p}+3)=1/9$. Put onto the scale of a single cluster, Chen et al., (2013) requires the faster rate of $\frac{1}{N}\lambda=O(N^{-8/9})$ than the $O(N^{-1/2})$ rate required by our theory in this special case. By this reasoning, condition (vi) is a weak assumption in this special case. Our conjecture is that this extends to more complicated settings although we omit such analysis in this work.
@@ B.1 Condition (vi)
$S^{-1/2}_{N}(\beta_{N})\left(\widehat{\beta}_{0,N}-\beta_{N}\right)$ =$~{}\overset{p=1}{=}~{}\frac{\widehat{\beta}^{(0)}_{N}-{\beta}^{(0)}_{N}}{\sqrt{S_{N}(\beta_{N}})}+\frac{{\beta}^{(0)}_{N}-{\beta}_{N}}{\sqrt{S_{N}(\beta_{N}})}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
[…]

### v2

@@ A.9 Initial pffr fit
where ${l}({Y}_{i,j}(s),\mathbb{X}_{i,j};\bm{\theta})$ is the log-likelihood evaluated on a single observation of the outcome ${Y}_{i,j}(s)\in\mathbb{R}$ from cluster $i$, at longitudinal observation $j$, at functional domain point $s$. We denote the model parameter $\bm{\theta}$, $\mathbb{S}$ as the penalty matrix, and $\Lambda_{0}$ as an associated diagonal matrix of smoothing parameters. As discussed in the main text, $\mathbb{X}_{i,j}\in\mathbb{R}^{p}$ is constructed as a product of pre-defined basis functions (e.g. B-splines) and the original covariates, $\bm{x}_{i,j}\in\mathbb{R}^{q}$. Model 8 is equivalent to adopting a correlation structure that assumes all observations $Y_{i,j}(s)$ are mutually independent across $i$, $j$, and $s$.
We fit model 8 with the refund package in R (Goldsmith et al., 2024) with the pffr function (Scheipl et al., 2016). This calls the mgcv package in R (Wood, 2017) to fit the model with the gam or bam functions. For example, for a model with two covariates, X1 and X2, we use the following code:
@@
where fam is the exponential dispersion family adopted to construct a pseudo-likelihood based estimating equation in the fGEE, spline.basis is the pre-specified basis spline from the mgcv family (e.g. bs, ps, tp), knots is the number of knots, and m.pffr is the penalty type. For example, as noted in the pffr function, bs.yindex = list(bs="ps", k=5, m=c(2, 1)) indicates 5 cubic B-splines bases with a first order difference penalty.
We use the above fit as both the initial estimate for the one-step fGEE and as benchmark method “pffr” in the simulations presented in Section 4 and throughout the simulations presented in the Appendix. We do not use functional random effects in any pffr fits.
A.10 Computational Details
Our implementation uses a number of R packages for estimation of nuisance parameters and to improve computational speed. We use the data.table package extensively to increase computational efficiency (Barrett et al., 2025). Our code structure was initially loosely based on the structure from the gee1step package that implements the one-step (non-functional) GEE (Lipsitz et al., 2017) available on the Github https://github.com/kgoldfeld/gee1step of Professor Keith Goldfeld. We use the Rfast package (Papadakis et al., 2018) to estimate the $\rho(s)$ for AR1 correlation and to speed up other standard computations. We use the MASS package to draw multivariate normals (Venables and Ripley, 2002). We use Allévius (2018) to estimate and invert covariance matrices with an AR1 structure when the time intervals are irregular. We use the SuperGauss package (Ling and Lysy, 2022) to invert AR1 matrices with regular time intervals. We use the sanic package to quickly invert other positive definite matrices (Kuschnig, 2023). As mentioned in the main text and previous Appendix section, we use mgcv (Wood, 2017) and refund (Goldsmith et al., 2024) packages to estimate initial fits and to (optionally) smooth the correlation parameters across the functional domain.
A.11 Description of the use of AI
We used Large Language Models (LLMs) to extend and optimize our existing package code (described below), modify our existing code to generate tables summarizing simulation results, and to generate ggplot2 code to generate Appendix Figure S1, which summarizes simulation results. Specifically, we used LLMs to write a computationally efficient version of the copula method to generate longitudinal functional outcomes given covariates that we used to simulate data. We did not use LLMs for other portions of our simulations. We extensively verified and tested code produced by LLMs with large simulation experiments. We used LLMs only in code added in revisions of an original manuscript, not in any code of the original manuscript. We did not use LLMs for writing or editing the manuscript, any of the mathematical proofs, or core methodological ideas.
The key areas we used LLMs for in extending our original package code are: 1) implementing an efficient data.table-based version of an initial Kronecker product working covariance; 2) the FPCA-based working covariance; 3) extension of the fGEE to link functions beyond log, logit and identity; 4) extension to estimating equations based on pseudo-distributions other than those deriving from Poisson, binomial, and Gaussian; 5) calculation of the small-$N$ effective degree of freedom inflation factor; 6) implementation of a computationally efficient version of the wild cluster bootstrap; and 7) implementation of speedups for the fast cluster CV (e.g., using Rcpp-based evaluations of the negative log-likelihood). We have acknowledged LLM use in the package documentation. When LLMs were used in generating code outside the package, we acknowledge their use in the R files (see Github repo: https://github.com/gloewing/fgee_onestep). All LLMs were accessed through the platforms provided to NIH Intramural Research Program by the Department of Health and Human Services (HHS), namely Gemini 2.5 Pro, Claude Sonnet 4.5 (while available to HHS employees), and ChatGPT 5.2 Pro.
@@ Appendix B Theory
B.1 Main Text Theorem Conditions
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
    $\bm{H}_{N}(\bm{\theta}_{N})=$ is invertible, $\mathbb{P}\left[\mathbb{P}_{N}\left(\nabla_{\bm{\theta}}\bm{U}_{N}(\mathbf{X}_{i},\bm{Y}_{i};\bm{\theta})\bigg|_{\bm{\theta}=\widehat{\bm{\theta}}^{(0)}_{N}}\right)~\text{non-singular}\right]=1$,
  5. (v)
  6. (vi)
Condition (i) is a mild smoothness condition that holds for all standard link functions (e.g. logit, log). Condition (ii) is also standard—we expect it holds across essentially all biomedical settings. Note that it could be replaced by weaker moments conditions on the estimating equation, and its derivatives. Condition (iii) is a sufficient condition for the estimating equation to be asymptotically normal and implies that $\bm{M}_{N}(\bm{\theta}_{N})$ is invertible for all $N\in\mathbb{N}$. Condition (iv) also states that the $\bm{H}_{N}(\bm{\theta}_{N})$, and its sample analogue, are invertible for all $N\in\mathbb{N}$. Condition (v) should hold when the limiting (unpenalized) estimating equation results in full rank limiting $\bm{M}_{N}$ and $\bm{H}_{N}$. This should hold when the design matrices, $\mathbb{X}_{i}$, are full rank. Finally condition (vi) is a statement about the rate of convergence of the initial estimator. In practice, when the $\widehat{\bm{\theta}}^{(0)}_{N}$ is estimated using a penalized unweighted estimating equation, this implies some conditions on the rates of convergence of the smoothing parameter values, $\Lambda_{0,N}$ and $\Lambda_{N}$. We provide an expanded discussion of this below.
B.2 General Result
Throughout, we fix an arbitrary matrix norm $\lVert\,\cdot\,\rVert$ (e.g., operator, Frobenius), and for a sequence of random matrices $(A_{N})_{N=1}^{\infty}$, we write $A_{N}=O_{\mathbb{P}}(1)$ if $\lVert A_{N}\rVert=O_{\mathbb{P}}(1)$, and $A_{N}=o_{\mathbb{P}}(1)$ if $\lVert A_{N}\rVert=o_{\mathbb{P}}(1)$. For symmetric matrix $A$, we write $\lambda_{\mathrm{min}}(A)$ and $\lambda_{\mathrm{max}}(A)$ for the smallest and largest eigenvalues of $A$, respectively. We begin by stating and proving a lemma that we use to prove our general theorem.
@@ B.2 General Result
$\{\bm{V}_{N}(\bm{\theta}_{N})\}^{-1/2}\left(\widehat{\bm{\theta}}_{N}^{*}-\bm{\theta}_{N}\right)\overset{d}{\to}\mathcal{N}(\bm{0}_{p},I_{p}).$
@@
    $\bm{H}_{N}(\bm{\theta}_{N})=$ is invertible and $\mathbb{P}\left[\mathbb{P}_{N}\left(\nabla_{\bm{\theta}}\bm{U}_{N}(Z;\bm{\theta})\bigg|_{\bm{\theta}=\widehat{\bm{\theta}}^{(0)}_{N}}\right)~\text{is non-singular}\right]=1$, for all $N\in\mathbb{N}$. Moreover, $\{\bm{H}_{N}(\bm{\theta}_{N})\}^{-1}=O(1)$ and $\left(\mathbb{P}_{N}\left[\nabla_{\bm{\theta}}\bm{U}_{N}(Z;\bm{\theta})\bigg|_{\bm{\theta}=\widehat{\bm{\theta}}^{(0)}_{N}}\right]\right)^{-1}=O_{\mathbb{P}}(1)$.
@@
Then the one-step estimator satisfies $\{\bm{V}_{N}(\bm{\theta}_{N})\}^{-1/2}\left(\widehat{\bm{\theta}}^{(1)}_{N}-\bm{\theta}_{N}\right)\overset{d}{\to}\mathcal{N}(\bm{0}_{p},{I}_{p})$.
@@
$\quad\quad+\left\{\mathbb{P}_{N}\left[\left.\nabla_{\bm{\theta}}\bm{U}_{N}(Z;\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{N}}\right]-\mathbb{E}\left(\left.\nabla_{\bm{\theta}}\bm{U}_{N}(Z;\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{N}}\right)\right\}.$
@@
B.3 Condition (vi)
To provide intuition for condition (vi) in the statement of Theorem B.2, we derive an interpretable set of conditions that imply condition (vi) in a special case of our general framework: a non-functional, univariate (i.e., $p=q=1$), non-clustered (i.e., $n_{i}=1~\forall~i\in[N]$) ridge regression that is weighted by the inverse of the working covariance matrix. This setting provides insight for the more general case: (1) the fully-iterated unweighted ridge is analogous to the initial penalized GEE estimator that is also unweighted (i.e., adopts an independence working covariance structure) and fully-iterated; (2) the fully-iterated weighted ridge is analogous to the fully-iterated and weighted (i.e., adopts some non-independence working covariance structure) penalized GEE. In this special case, we show that if the scaled smoothing parameters for the initial estimate satisfy $\frac{1}{N}\lambda_{0,N}=O(N^{-1/2})$, and for the weighted estimator satisfy $\frac{1}{N}\lambda_{N}=O(N^{-1/2})$, then condition (vi) holds. That is, if each of these scaled smoothing parameters individually (i.e., no conditions are required jointly on these rates) go to zero fast enough, then $\sqrt{N}\{\bm{M}_{N}(\bm{\theta}_{N})\}^{-1/2}\bm{H}_{N}(\bm{\theta}_{N})(\widehat{\bm{\theta}}^{(0)}_{N}-\bm{\theta}_{N})=O_{\mathbb{P}}(1)$. Importantly, the smoothing parameter rates we require are weaker than (i.e., implied by) the rates needed for the theoretical properties described in Chen et al. (2013). For example, even for the small knot setting, Chen et al. (2013) require that $\lambda=O(N^{\gamma})$ for $\gamma\leq(\tilde{p}+2-\tilde{q})/(2\tilde{p}+3)$. This implies, for instance, that for a $\tilde{p}^{th}$ order truncated polynomial using cubic B-splines, $\gamma\leq(\tilde{p}+2-(\tilde{p}+1))/(2\tilde{p}+3)=1/9$. Put onto the scale of a single cluster, Chen et al. (2013) requires the faster rate of $\frac{1}{N}\lambda=O(N^{-8/9})$ than the $O(N^{-1/2})$ rate required by our theory in this special case. By this reasoning, condition (vi) is a weak assumption in this special case. Our conjecture is that this extends to more complicated settings although we omit such analysis in this work.
@@ B.3 Condition (vi)
$S^{-1/2}_{N}(\beta_{N})\left(\widehat{\beta}_{0,N}-\beta_{N}\right)$ $~\overset{p=1}{=}~\frac{\widehat{\beta}^{(0)}_{N}-{\beta}^{(0)}_{N}}{\sqrt{S_{N}(\beta_{N}})}+\frac{{\beta}^{(0)}_{N}-{\beta}_{N}}{\sqrt{S_{N}(\beta_{N}})}$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
$~~=O(1)*\underbrace{\frac{\widehat{\beta}^{(0)}_{N}-{\beta}^{(0)}_{N}}{\sqrt{S_{N}^{(0)}(\beta_{N}^{(0)}})}}_{\begin{subarray}{c}\overset{d}{\to}N(0,1)\\ \text{under weak}\\ \text{CLT conditions}\end{subarray}}~+~A$